OpenStudy (anonymous):

The sum of twice a number and thirteen is fifty-five. Find the number. Help guys

OpenStudy (calculusxy):

2x + 13 = 55 Solve for x.

OpenStudy (calculusxy):

Do you know how to do this?

OpenStudy (anonymous):

2x=55-13 2x/2=42/2 x=21

OpenStudy (calculusxy):

Let's check. 2(21) + 13 42 + 13 55 Great job!